OpenStudy (idku):

weird but no one taught me in school how to take a derivative of an absolute value... tell me is I am doing it correctly.

OpenStudy (idku):

I am taking the derivative of\[\left| x^3 \right|\] re-writing, \[\sqrt{(x^3)^2}\] I will post this so far, because the post button is hiding

OpenStudy (idku):

I mean the equation editor is hiding, anyway...

myininaya (myininaya):

You could always write it as a piecewise function and differentiate OR use the chain rule and use the following: \[(|x|)'=\frac{x}{|x|}\]

OpenStudy (idku):

\[\frac{x}{dx}~\sqrt{(x^3)^2}~=~\frac{1}{2\sqrt{(x^3)^2}} \times 6x^5\]so, I would be getting: \[\frac{x}{dx}~\sqrt{(x^3)^2}~= \frac{3x^5}{\left| x^3 \right|}\]

OpenStudy (idku):

am I differentiating correctly?

OpenStudy (idku):

if so, then also, I should be able to take out the x^2, or not?

myininaya (myininaya):

Looks good. \[(|x^3|)'=\frac{x^3}{|x^3|}(x^3)'=\frac{x^3}{|x^3|}3x^2=\frac{3x^5}{|x^3|}\]

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

a bit complicated :| when u deal with absolute value remember this REWRITE IT AS PIECEWISE

OpenStudy (idku):

can I make it \[\frac{3x^5}{\left| x^3 \right|}=\frac{3x^3}{\left| x\right|}\] ?

myininaya (myininaya):

x^2>0 so |x^3|=|x^2*x|=|x^2|*|x|=x^2|x|

myininaya (myininaya):

x^2>=0*

OpenStudy (idku):

yes, I know that;)

OpenStudy (idku):

so I took x^2 out correctly?

myininaya (myininaya):

yes x^2/x^2=1

OpenStudy (idku):

I mean without dealing with imaginary values.

OpenStudy (idku):

because then I wouldn't, because |-i| doesn't equal -i

OpenStudy (idku):

but for real numbers, I took out the x^2 from top and bottom correctly... I suppose.

OpenStudy (michele_laino):

I think that since I can re-write your function as below: f(x)= x^3 if x>=0 and f(x)=-x^, x<0, then the derivative is: